OpenStudy (anonymous):

3/7+2/7 1/9+4/9+2/9 3/5-2/5

OpenStudy (anonymous):

i will become ur fan

OpenStudy (anonymous):

no joke plzzzzzzzzzzzzzzzzzzzzzzz

OpenStudy (anonymous):

3/7 + 2/7 = 5/7 1/9 + 4/9 + 2/9 = 7/9 3/5-2/5= 1/5 When they have a common number at the bottom, you just add the top part of the fraction

OpenStudy (anonymous):

thx